# Rapid Word Learning Through Meta In-Context Learning

### v1

@@ 1 Introduction
Children can quickly learn a new word, or at least make meaningful inferences about its meaning, given only a few examples of its usage (Carey and Bartlett, 1978; Bloom, 2000). For example, suppose a child who did not know the word ski hears the following mentions of the word (without visual examples): “Susie learned to ski last winter”, “People ski on tall mountains where there’s lots of snow”, and “I saw Susie ski fast down the snowy mountain.” From these usage examples, the child might infer that ski is a verb for a winter activity involving sliding down snowy mountains, and could begin understanding and using the word appropriately in new contexts. This ability to generalize and use a new word in novel contexts from just a few examples reflects children’s remarkable data efficiency in language learning, allowing them to quickly acquire vocabulary without requiring tens or hundreds of examples per word.
@@
To demonstrate the data efficiency of our method, we train language models from scratch with Minnow using small datasets: a corpus of child-directed speech (CHILDES; MacWhinney, 1992) and a corpus approximating the word count a child encounters during language acquisition (BabyLM-10M; Warstadt et al., 2023). To foreshadow our results, we find that our method’s performance on few-shot classification of new words from these datasets approaches that of the pre-trained Llama-3 8B (Llama Team, Meta AI, 2024), which was trained on vastly more data. This highlights how this ability can be developed from human-scale child-input data rather than the orders-of-magnitude larger datasets typically used to train LLMs.
@@ 3 Method
The goal of our method, Minnow, is to enable a model to infer the meaning of a new word from a few examples of its usage so it can understand and generate novel usage examples of the word, coherently and systematically combining it with other words in new contexts. To achieve this, Minnow trains the model to generate another usage example of the new word—a task that, when sufficiently challenging, requires mastery of this ability. Minnow is a general framework that can be applied to both training a model from scratch and finetuning a pre-trained model. After describing the method, we introduce the training data we use, a held-out word classification task for model evaluation and hyperparameter tuning, and how we use the off-the-shelf Llama-3 8B as a baseline for our experiments.
@@ 3.1 Method: Minnow
When we train a model from scratch, we also provide episodes of language modeling (without masked new tokens) to further facilitate language learning, as illustrated in Figure 1 (bottom). Each of these episodes consists of the same number of $K$ randomly sampled unrelated sentences, without new words. We concatenate them in the same format and train the model to perform next-token prediction on the concatenated sequences. Training batches of language modeling episodes interleave with the batches of meta-learning episodes. The model can determine whether an episode is for meta-learning or language modeling from whether the special placeholder token occurs in the first sentence.
@@ 3.3 Held-out Word Classification
Given a query example sentence $q$ that uses a new word and a set of $C$ candidate words $\{w^{(c)}\}_{c=1}^{C}$, the task for the model is to match the query example to the most suitable one among the $C$ candidate words. Each $w^{(c)}$ is represented by a context containing a set of $K-1$ study examples $\{x^{(c)}_{k}\}_{k=1}^{K-1}$ illustrating its usage. The context of $w^{(c)}$ is a sequence in the same format as the first $K-1$ examples in a training episode, ending with a separator token (e.g., <sep>): <sep> $x^{(c)}_{1}$ <sep> $\cdots$ <sep> $x^{(c)}_{K-1}$ <sep>. The query example is formatted as a continuation sequence of the context: $q$ <sep>. This formatting ensures that concatenating a context sequence and a query sequence results in a sequence with $K$ examples, just like a sequence for a meta-learning training episode. To determine the best match, we compute the conditional likelihood of the query sequence given the context: $p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The model predicts the word corresponding to the context with the highest likelihood: $\operatorname*{arg\,max}_{c}p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The prediction is correct if it is the ground-truth word in the query $q$.
We evaluate each model (trained from scratch or finetuned) by measuring the classification accuracy on held-out meta-learned words from the validation or test portions of the model’s training or finetuning corpus. For each evaluation, we group $C$ distinct meta-learned words into a $C$-way classification task. For each word, we sample $K-1$ study examples and one query example to construct the task. See Appendix C for additional details on task construction.
3.4 Baseline: Off-the-shelf Llama-3 8B
@@ Results
Models trained from scratch on $K=5$ examples per episode sampled from CHILDES and BabyLM-10M achieve test accuracies of 72% and 77%, respectively, on the 4-way ($C=4$) classification task. These results are substantially higher than random chance (25%) and close to the 71% and 78% accuracies achieved by Llama-3 8B baseline, which was pre-trained on orders of magnitude more data. We provide results in additional settings, including experiments with $K=10$ examples on CHILDES and 8-way ($C=8$) classification, in Appendix C, Table 6. Across all settings, models trained from scratch consistently achieve accuracies well above chance and within a 3% margin of the Llama-3 8B baseline. These findings (on CHILDES in particular) demonstrate that few-shot word learning can be effectively acquired using our method, even with human-scale child-input data.
@@ 5 Finetuning Pre-trained LLMs
In this section, we test if our method can improve pre-trained LLMs’ in-context few-shot word learning abilities. We finetune Llama-3 8B with Minnow three times on the meta-learning component of BabyLM-10M, each run with $K=5$ examples per episode and a different random seed. [^5] We focus on finetuning models on BabyLM-10M in this section, since it is more diversified and usually yields better results than CHILDES. We refer to the models finetuned with Minnow as Minnow models. We do not include the language modeling components since the LLM already learned a large vocabulary and is capable of language modeling. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, same as for the baseline (Section 3.4). We freeze all of the model’s parameters except the input and output embeddings of these two special tokens. We initialize the embeddings of these two special tokens as the mean of all other input/output embeddings (Hewitt, 2021). We select the checkpoint for each run and tune the learning rate in the same way as when training from scratch, except that we do not apply weight decay (Section 4). See Appendix B for more details on text formatting, tokenization, and training hyperparameters including batch size and learning rate (with scheduling). In the following, we evaluate the Minnow models and baselines on a series of tasks.
@@ 5.1 Held-out Word Classification
We first evaluate models on the held-out word classification task (Section 3.3). Finetuning Llama-3 8B with Minnow boosts the test 4-way ($C=4$) classification accuracy from the baseline level of 78% to 87% on BabyLM-10M (and from 71% to 79% on CHILDES). We provide results for additional values of $K$ and $C$ in Appendix C, Table 6; broadly, across all settings, the Minnow model improves test accuracy by 8–10% over the Llama-3 8B baseline. These findings show that Minnow finetuning effectively improves the pre-trained LLM’s in-context few-shot word learning ability.
@@ 5.2 Syntactic Category Classification
Note that, unlike the previous task, query examples are semantically unrelated to the study examples in this task, thus excluding the shortcut of semantic coherence. Below, we report the mean accuracies across the three runs.
@@ Results
We first find that the Llama-3 8B baseline achieves 64% accuracy on this task, which is higher than random chance (50%), suggesting that it can infer the syntactic categories of new words in one shot and generalize them to novel contexts. The Minnow model improves accuracy to 83%, a 19% increase over the baseline. Fine-grained results from models finetuned with Minnow (and trained from scratch) are provided in Appendix D. We find in all settings that the Minnow model improves accuracy by 11–26% compared to the baseline on all pairs of categories. These improvements show that Minnow finetuning effectively helps in learning the syntactic categories of new words and generalizing accordingly. In addition, note that our models are not specifically finetuned on this syntactic category classification task and dataset, demonstrating the generality of the acquired word learning ability.
@@ 5.3 New Usage Example Generation
The two tests we have described so far evaluate models in a discriminative setting. Here, we quantitatively and qualitatively evaluate if models use the new word appropriately in a generative setting. For a Minnow model finetuned with $K$ examples per episode, we evaluate it by showing it $K-1$ in-context study examples, formatted as a sequence in the classification setting (Section 3.3). We ask the model to do what it was trained for: We prompt the model with this sequence of study examples, and because the sequence ends with a separator token, the model will continue the sequence by generating a new usage example, ending with another separator token as End-Of-Sequence.
We sample study examples from two datasets: the BabyLM-10M test portion in Section 3.2 and the Chimera dataset (Lazaridou et al., 2017). The Chimera dataset was specifically constructed for few-shot word learning. It has 33 different new words for learning, each referring to a “chimera” concept, i.e., a mixture of two existing and related concepts (e.g., a cello and a bagpipe). The usage examples of a new word are sentences using one of the components of the chimera, randomly extracted from a large corpus. See Appendix F for additional details of the dataset and our preprocessing.
For the quantitative evaluation, we compare a pair of new usage examples generated from Llama-3 8B baseline and a Minnow model finetuned from it. The comparison is simulated as a head-to-head competition following the methodology in the definition generation section of Teehan et al. (2024). Specifically, we provide GPT-4o (OpenAI, 2024) the same $K-1$ study examples in a list format with a pseudo-word “dax” as the placeholder for the word, as in the baseline (without the last separator; Section 3.4), followed by a question “Which of the following is a better next example for the word ‘dax’, or they tie?” with three shuffled options, including the two generations and one “Tie”. (See Appendix E for detailed settings of prompting.) The choice of GPT-4o decides whether and which one model wins the competition, or whether the models were tied in quality. For the qualitative evaluation, we manually pick meta-learned words (shown in Table 2 and Tables 1011, and 12 in Appendix F) and examine the syntactic correctness and semantic appropriateness of the generated examples.
@@
instruction-tuned baseline 41 52 33
@@
Table 1: Percentages of wins of each model when comparing the generations from Llama-3 8B baseline (pre-trained to instruction-tuned) with a Minnow model finetuned from that baseline, judged by GPT-4o. The left two datasets are for new usage example generation in Section 5.3, and the right-most one is for definition generation in Section 5.4. Each new example or definition is generated by greedy decoding. (Results of top-p sampled generations are shown in Table 9 in Appendix E.) The percentage of ties is the remaining after subtracting the win percentages of the two models. GPT-4o more frequently chooses the Minnow model as the winner compared to the corresponding baseline in all settings except for the instruction-tuned variant on Chimera.
@@ Results
For the quantitative evaluation, Table 1 shows the percentages of wins of each of the baseline and the Minnow model on both the BabyLM-10M test portion and Chimera. Across all settings, the Minnow model wins more often than the corresponding baseline except for the instruction-tuned variant on Chimera, demonstrating the improvement brought by Minnow. For the qualitative evaluation, Table 2 shows a word picked from the BabyLM-10M test portion along with its study and generated examples. See Appendix F for additional examples from the BabyLM-10M test portion (Tables 10 and 12) and Chimera (Table 11) and detailed analysis of both the baseline and the Minnow model’s generations. A manual analysis of these generated examples reveals that the Minnow model more often generates syntactically correct and semantically plausible new usage examples compared to the baseline, confirming that Minnow finetuning improves the ability to understand and use a new word. Nevertheless, in several cases, the Minnow model still shows obvious syntactic and factual errors and merely rewords the study examples.
@@ 5.4 Definition Generation
To further probe how well Minnow finetuning helps the model understand a new word, we prompt each model to generate a definition for the word given one or a few usage examples. We again use the methodology of Teehan et al. (2024) for definition generation and evaluation, as well as the two evaluation datasets they used: CoLLEGe-DefGen, which they created, and the Oxford dataset (Gadetsky et al., 2018). CoLLEGe-DefGen was constructed by selecting 954 words from WordNet (Miller, 1995) and prompting GPT-4 (OpenAI, 2023) to generate one definition and five usage examples for each word. The model generates a definition from one, two, or three usage examples sampled for each word in this dataset (i.e., in 1-, 2-, or 3-shot settings). The Oxford test set consists of 12,232 words, each with a definition and a usage example collected from the Oxford Dictionary. The model generates a definition from the only usage example for each word in this dataset (i.e., in a 1-shot setting). To generate a definition, we prompt the model with the sequence of the usage example(s) (as in Section 5.3) followed by “The word [new-token] in the above sentence(s) is defined as "” [^6] The prompt ends with a double quotation mark (”), so that the model will continue with a definition ending at another double quotation mark. This makes extracting definitions easy. ([new-token] is instead the placeholder token or pseudoword, as appropriate). For additional comparisons with models expected to do especially well on this task, we also evaluate specialized definition-generation models: Giulianelli et al.’s (2023) FLAN-T5 models (Chung et al., 2024). See Appendix G for details of data preprocessing and the specialized models.
For the quantitative evaluation, we perform two types of comparison. The first type compares the model-generated and ground-truth definitions for each word by computing BERTScore F1 (Zhang et al., 2020) and ROUGE-L (Lin, 2004). The second type compares a pair of definitions generated from Llama-3 8B baseline and a Minnow model finetuned from it. Similarly to what we did in Section 5.3, we ask GPT-4o a question (without usage examples): “Which of the following is a better definition for the word ‘Word’, or they tie?” where Word is the ground-truth word form, followed by three shuffled options including the two generated definitions and one “Tie” (see Appendix E for detailed prompting settings). [^7] We only perform this comparison on the CoLLEGe-DefGen dataset due to the large scale of the Oxford dataset. For the qualitative evaluation, we manually inspect 1-shot generated definitions for words from each dataset (presented in Table 4 and Tables 15 and 16 in Appendix G).
@@
instruction-tuned baseline 85.3 17.6 83.6 12.5
@@
Table 3: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions. See Table 13 in Appendix G for results from all models. We generate a definition from only one example (1-shot). All definitions are generated with greedy decoding. Scores of Minnow models are averaged across three runs. Finetuning with Minnow improves the baseline models on both datasets and both metrics, and the Minnow model finetuned from the instruction-tuned variant of Llama-3 8B performs the best.
@@
After his thorough inspection of the antique pocket watch, the bespectacled collector sighed, claiming it was a [new-token], much to the seller’s disappointment. a thing that is not genuine or authentic a deception or trick swiz
@@
Table 4: Definitions for two words from CoLLEGe-DefGen generated by the Minnow model finetuned from instruction-tuned Llama-3 8B with greedy decoding. Each definition is generated using the single example sentence shown and provided in context. The generated definitions managed to infer the core semantic features from the examples, though they are not precise enough compared to the true definitions. In this first example, the Minnow definition for the word “swiz” captures the word’s core meaning of fakeness, which is a reasonable inference from the example, but misses the intentional aspect, a nuance of the true definition. In the second example, the Minnow definition for “categorical imperative” captures the core meaning of obligation, which is a reasonable contrast to the businessman’s greed, but misses the “unconditionally and universally binding” aspect in the true definition.

### v2

@@ 1 Introduction
Children can quickly learn a new word, or at least make meaningful inferences about its meaning, given only a few examples of its usage (Carey and Bartlett, 1978; Bloom, 2000). For example, suppose a child who did not know the word ski hears the following mentions of the word (without visual examples): “Susie learned to ski last winter”, “People ski on tall mountains where there’s lots of snow”, and “I saw Susie ski fast down the snowy mountain.” From these usage examples, the child might infer that ski is a verb for a winter activity involving sliding down snowy mountains, and could begin understanding and using the word appropriately in new contexts. [^1] Learning a new word is often equivalent to learning a new concept (Murphy, 2002). Therefore, we equate word learning to concept learning throughout the paper. This ability to generalize and use a new word in novel contexts from just a few examples reflects children’s remarkable data efficiency in language learning, allowing them to quickly acquire vocabulary without requiring tens or hundreds of examples per word.
@@
To demonstrate the data efficiency of our method, we train language models from scratch with Minnow using small datasets: a corpus of child-directed speech (CHILDES; MacWhinney, 1992) and a corpus approximating the word count a child encounters during language acquisition (BabyLM-10M; Warstadt et al., 2023). To foreshadow our results, we find that our method’s performance on few-shot classification of new words from these datasets approaches that of the pre-trained Llama-3 8B (Meta AI, 2024), which was trained on vastly more data. This highlights how this ability can be developed from human-scale child-input data rather than the orders-of-magnitude larger datasets typically used to train LLMs.
@@ 3 Method
The goal of our method, Minnow, is to enable a model to infer the meaning of a new word from a few examples of its usage so it can understand and generate novel usage examples of the word, coherently and systematically combining it with other words in new contexts. To achieve this, Minnow trains the model to generate another usage example of the new word—a task that, when sufficiently challenging, requires mastery of this ability. Minnow is a general framework that can be applied to both training a model from scratch and finetuning a pre-trained model. After describing the method, we introduce the training data we use, a held-out word classification task for model evaluation and hyperparameter tuning, and how we use the off-the-shelf Llama model and the CoLLEGe model (introduced in Section 2.2) as baselines for our experiments.
@@ 3.1 Method: Minnow
When we train a model from scratch, we also provide episodes of language modeling (without placeholder tokens) to further facilitate language learning, as illustrated in Figure 1 (bottom). Each of these episodes consists of the same number of $K$ randomly sampled unrelated sentences, without new words. We concatenate them in the same format and train the model to perform next-token prediction on the concatenated sequences. Training batches of language modeling episodes interleave with the batches of meta-learning episodes. The model can determine whether an episode is for meta-learning or language modeling from whether the special placeholder token occurs in the first sentence.
@@ 3.3 Held-out Word Classification
Given a query example sentence $q$ that uses a new word and a set of $C$ candidate words $\{w^{(c)}\}_{c=1}^{C}$, the task for the model is to match the query example to the most suitable one among the $C$ candidate words. Each $w^{(c)}$ is represented by a context containing a set of $K-1$ study examples $\{x^{(c)}_{k}\}_{k=1}^{K-1}$ illustrating its usage. (Note that in all query and study examples, the occurrences of the new word are replaced with the same special placeholder token, e.g., [new-token], as described in Section 3.1.) The context of $w^{(c)}$ is a sequence in the same format as the first $K-1$ examples in a training episode, ending with a separator token (e.g., <sep>): <sep> $x^{(c)}_{1}$ <sep> $\cdots$ <sep> $x^{(c)}_{K-1}$ <sep>. The query example is formatted as a continuation sequence of the context: $q$ <sep>. This formatting ensures that concatenating a context sequence and a query sequence results in a sequence with $K$ examples, just like a sequence for a meta-learning training episode. To determine the best match, we compute the conditional likelihood of the query sequence given the context: $p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The model predicts the word corresponding to the context with the highest likelihood: $\operatorname*{arg\,max}_{c}p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$. The prediction is correct if it is the ground-truth word in the query $q$.
We evaluate each model (trained from scratch or finetuned) by measuring the classification accuracy on held-out meta-learned words from the validation or test portions of the model’s training or finetuning corpus. For each evaluation, we group $C$ distinct meta-learned words into a $C$-way classification task. For each word, we sample $K-1$ study examples and one query example to construct the task. See Appendix C for an example task and additional details on task construction.
3.4 Baselines
3.4.1 Off-the-shelf Llama model
@@
3.4.2 CoLLEGe
An alternative to Minnow is to generate new embeddings for new words (Section 2.2). For instance, CoLLEGe uses meta-learning to train an additional transformer encoder layer over a pre-trained MLM, RoBERTa (Liu et al., 2019), to generate new input and output embeddings for a new word token (e.g., [new-token]) based on a set of study examples. The new token is then used by the pre-trained Llama-2 7B (Touvron et al., 2023). We use the original checkpoint of CoLLEGe as another baseline. In the held-out word classification task (Section 3.3), the conditional likelihood $p_{\textrm{LM}}(q\mid x^{(c)}_{1},\ldots,x^{(c)}_{K-1})$ is computed by using only the input and output embeddings generated by CoLLEGe based on the study examples $x^{(c)}_{1},\ldots,x^{(c)}_{K-1}$. For fair comparison between Minnow and CoLLEGe, we also finetuned from Llama-2 7B with Minnow and compare this Minnow model to the CoLLEGe model. We also run off-the-shelf Llama-2 7B as an additional baseline, which is the same as described in Section 3.4.1.
@@ Results
Models trained from scratch on $K=5$ examples per episode sampled from CHILDES and BabyLM-10M achieve test accuracies of 72% and 77%, respectively, on the 4-way ($C=4$) classification task. These results are substantially higher than random chance (25%) and close to the 71% and 78% accuracies achieved by Llama-3 8B baseline (70% and 78% accuracies by Llama-2 7B baseline), which was pre-trained on orders of magnitude more data. We provide results in additional settings, including experiments with $K=10$ examples on CHILDES and 8-way ($C=8$) classification, in Appendix C, Table 6. Across all settings, models trained from scratch consistently achieve accuracies well above chance and within a 3% margin of the Llama-3 8B baseline. These findings (on CHILDES in particular) demonstrate that few-shot word learning can be effectively acquired using our method, even with human-scale child-input data.
@@ 5 Finetuning Pre-trained LLMs
In this section, we test if our method can improve pre-trained LLMs’ in-context few-shot word learning abilities. We finetune Llama-3 8B (and Llama-2 7B) with Minnow three times on the meta-learning component of BabyLM-10M, each run with $K=5$ examples per episode and a different random seed. [^6] We focus on finetuning models on BabyLM-10M in this section, since it is more diversified and usually yields better results than CHILDES. We refer to the models finetuned with Minnow as Minnow models. We do not include the language modeling components since the LLM already learned a large vocabulary and is capable of language modeling. We finetune from both the pre-trained and instruction-tuned variants of Llama-3 8B, but we refer to the models finetuned from the pre-trained variant by default, same as for the off-the-shelf baseline (Section 3.4.1). We freeze all of the model’s parameters except the input and output embeddings of these two special tokens. [^7] See Appendix I for its effect on other general capabilities. We initialize the embeddings of these two special tokens as the mean of all other input/output embeddings (Hewitt, 2021). We select the checkpoint for each run and tune the learning rate in the same way as when training from scratch, except that we do not apply weight decay (Section 4). See Appendix B for more details on text formatting, tokenization, and training hyperparameters including batch size and learning rate (with scheduling). In the following, we evaluate the Minnow models and baselines on a series of tasks.
@@ 5.1 Held-out Word Classification
We first evaluate models on the held-out word classification task (Section 3.3). Finetuning Llama-3 8B with Minnow boosts the test 4-way ($C=4$) classification accuracy from the baseline level of 78% to 87% on BabyLM-10M (and from 71% to 79% on CHILDES). We provide results for additional values of $K$ and $C$ and for models based on Llama-2 7B (including CoLLEGe) in Appendix C, Table 6; broadly, across all settings, the Minnow model improves test accuracy by 8–10% over the Llama-3 8B baseline, while the CoLLEGe model performs 3–15% worse than the Llama-2 7B baseline. These findings show that Minnow finetuning effectively improves the pre-trained LLM’s in-context few-shot word learning ability.
@@ 5.2 Syntactic Category Classification
Note that, unlike the previous task, query examples are semantically unrelated to the study examples in this task, thus excluding the shortcut of semantic coherence. Below, we report the mean accuracies across three runs.
@@ Results
We first find that the Llama-3 8B baseline achieves 66% accuracy on this task, which is higher than random chance (50%), suggesting that it can infer the syntactic categories of new words in one shot and generalize them to novel contexts. The Minnow model improves Llama-3 8B’s accuracy to 83%, a 17% increase over the baseline. Meanwhile, both Minnow and CoLLEGe improve Llama-2 7B’s accuracy from 69% to 80%, which is an 11% increase. Fine-grained results from models finetuned with Minnow, trained from scratch, and CoLLEGe are provided in Appendix D. We find in all settings that the Minnow model improves accuracy by 4–23% compared to the baseline on all pairs of categories. These results show that Minnow finetuning effectively helps in learning the syntactic categories of new words and generalizing accordingly, and is comparable to CoLLEGe in improvements. In addition, note that our models are not specifically finetuned on this syntactic category classification task and dataset, demonstrating the generality of the acquired word learning ability.
@@ 5.3 New Usage Example Generation
Llama-3 8B Instruct baseline 41 52 33
@@
Llama-2 7B +CoLLEGe 13 33 4
+Minnow 74 58 47
Table 1: Percentages of wins of each model when comparing the generations from the pairs of models in each box, judged by GPT-4o. In the top two ruled rows, we compare Llama-3 8B baseline (pre-trained to instruction-tuned) with a Minnow model finetuned from that baseline. In the bottom-most ruled row, we compare CoLLEGe with the Minnow model finetuned from the pre-trained Llama-2 7B. Each new example or definition is generated by greedy decoding. (Results of top-p sampled generations are shown in Table 10 in Appendix E.) The percentage of ties is the remaining after subtracting the win percentages of the two models. GPT-4o more frequently chooses the Minnow model as the winner compared to the corresponding baseline model in all settings except for the instruction-tuned variant on Chimera.
@@
The two tests we have described so far evaluate models in a discriminative setting. Here, we quantitatively and qualitatively evaluate if models use the new word appropriately in a generative setting. For a Minnow model finetuned with $K$ examples per episode, we evaluate it by showing it $K-1$ in-context study examples, formatted as a sequence in the classification setting (Section 3.3). We ask the model to do what it was trained for: We prompt the model with this sequence of study examples, which ends with a separator token, so the model will continue the sequence by generating a new usage example, ending with another separator token as End-Of-Sequence. For CoLLEGe, we generate a new example using the prompt “A single example sentence using the word ‘[new-token]’ (in one line):”.
We sample study examples from two datasets: the BabyLM-10M test portion in Section 3.2 and the Chimera dataset (Lazaridou et al., 2017). The Chimera dataset was specifically constructed for few-shot word learning. It has 33 different new words for learning, each referring to a “chimera” concept, i.e., a mixture of two existing and related concepts (e.g., cello and bagpipe). See Appendix F for more details of the dataset.
For the quantitative evaluation, we compare a pair of new usage examples generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. The comparison is simulated as a head-to-head competition following Teehan et al. (2024). Specifically, we provide GPT-4o (OpenAI, 2024) the same $K-1$ study examples in a list format with a pseudo-word “dax” as the placeholder for the word, as in the off-the-shelf baseline (without the last separator; Section 3.4.1), followed by a question “Which of the following is a better next example for the word ‘dax’, or they tie?” with three shuffled options, including the two generations and one “Tie”. (See Appendix E for prompting details.) The choice of GPT-4o decides whether and which one model wins the competition, or whether the models were tied in quality. For the qualitative evaluation, we manually pick meta-learned words (Table 2 and Appendix F) and examine the syntactic correctness and semantic appropriateness of the generated examples.
@@ Results
For the quantitative evaluation, Table 1 shows the percentages of wins of each of the baseline and the Minnow model on both the BabyLM-10M test portion and Chimera. Across all settings, the Minnow model wins more often than the corresponding baseline except for the instruction-tuned variant on Chimera, demonstrating the improvement brought by Minnow and its better performance compared to CoLLEGe. For the qualitative evaluation, Table 2 shows a word picked from the BabyLM-10M test portion along with its study and generated examples. See Appendix F for additional examples from the BabyLM-10M test portion and Chimera and detailed analysis of the generations. A manual analysis of these generated examples reveals that the Minnow model more often generates syntactically correct and semantically plausible new usage examples compared to the baseline, confirming that Minnow finetuning improves the ability to understand and use a new word. Nevertheless, in several cases, the Minnow model still shows obvious syntactic and factual errors and merely rewords the study examples.
@@ 5.4 Definition Generation
Llama-3 8B Instruct baseline 85.3 17.6 83.6 12.5
@@
Llama-2 7B +CoLLEGe 84.0 16.3 83.3 14.1
+Minnow 82.9 18.0 83.6 15.6
FLAN-T5 XL +DefInstr baseline 83.1 12.4 84.9 19.4
Table 3: Quantitative evaluation of 1-shot generated definitions by comparing them with ground-truth definitions. We sample an example per word from CoLLEGe-DefGen. All definitions are generated with greedy decoding. “FLAN-T5 XL +DefInstr” is a specialized definition-generation model from Giulianelli et al. (2023). “baseline” means using a pseudo-word ‘wug’ as the placeholder. Scores of Minnow models (“+Minnow”) are averaged across 3 runs. Finetuning Llama-3 8B with Minnow improves the baseline models on both datasets and both metrics, and the Minnow model finetuned from the instruction-tuned variant of Llama-3 8B performs the best. The instruction-tuned variant of Llama-3 8B (“Llama-3 8B Instruct”) is better than the pre-trained variant (“Llama-3 8B”) on definition generation likely due to its better instruction-following ability. The Minnow model also beats CoLLEGe except for the BERTScore F1 on CoLLEGe-DefGen. The specialized model (“FLAN-T5 XL +DefInstr”) performs the best on Oxford, but note that it is a model specialized in definition generation and was finetuned on Oxford.
@@
Table 4: Definition for a word from CoLLEGe-DefGen generated by the Minnow model finetuned from instruction-tuned Llama-3 8B with greedy decoding. The definition is generated using the single example sentence shown and provided in context. The generated definition managed to infer the core semantic features from the examples, though they are not precise enough compared to the true definitions. In the example, the Minnow definition for “categorical imperative” captures the core meaning of obligation, which is a reasonable contrast to the businessman’s greed, but misses the “unconditionally and universally binding” aspect in the true definition.
To further probe how well Minnow finetuning helps the model understand a new word, we prompt each model to generate a definition for the word given one or a few usage examples. We again follow Teehan et al. (2024) for definition generation and evaluation, as well as the two evaluation datasets they used: CoLLEGe-DefGen, which they created, and the Oxford dataset (Gadetsky et al., 2018). CoLLEGe-DefGen was constructed by selecting 954 words from WordNet (Miller, 1995) and prompting GPT-4 (OpenAI, 2023) to generate one definition and five usage examples for each word. The model generates a definition from one, two, or three usage examples sampled for each word in this dataset (i.e., 1-, 2-, or 3-shot). The Oxford test set consists of 12,232 words, each with a definition and a usage example collected from the Oxford Dictionary. The model generates a definition from the only usage example for each word in this dataset (i.e., 1-shot). To generate a definition, we prompt Llama and Minnow models with the sequence of the usage example(s) (as in Section 5.3) followed by “The word [new-token] in the above sentence(s) is defined as "” [^8] The prompt ends with a double quotation mark (”), so that the model will continue with a definition ending at another double quotation mark. This makes extracting definition easy. ([new-token] is instead the placeholder token or pseudoword, as appropriate). For CoLLEGe, we use the same prompt but without the in-context usage example(s). See Appendix G for details of data preprocessing and additional specialized definition-generation models from comparison (Giulianelli et al., 2023).
For the quantitative evaluation, we perform two types of comparison. The first type compares the model-generated and ground-truth definitions for each word by computing BERTScore F1 (Zhang et al., 2020) and ROUGE-L (Lin, 2004). The second type compares a pair of definitions generated from Llama-3 8B baseline and a Minnow model finetuned from it, or the CoLLEGe baseline and a Minnow model finetuned from Llama-2 7B. Similarly to what we did in Section 5.3, we ask GPT-4o a question (without usage examples): “Which of the following is a better definition for the word ‘Word’, or they tie?” where Word is the ground-truth word form, followed by three shuffled options including the two generated definitions and one “Tie” (see Appendix E for detailed prompting settings). [^9] We only perform this comparison on the CoLLEGe-DefGen dataset due to the large scale of the Oxford dataset. For the qualitative evaluation, we manually inspect 1-shot generated definitions for words from each dataset (presented in Table 4 and Tables 18 and 19 in Appendix G).